Which expression is equivalent to the one below?

(2^6 x 9^6) divided 18^{14}
1
Expert's answer
2019-10-06T07:02:55-0400

Solution:


We need to find the Equivalent Expression to ( 26 x 96 ) Divided by 1814


Step1:

In the Multiplication of two terms with same exponent, we can multiply the bases and write the same exponent for the result

of multiplication.


For example,

am x b m = ( a x b )m


so, 26 x 96 = ( 2 x 9 )6 = 186


Step 2:

In the division of two terms with same base and different exponents, we can write the base at the numerator and the exponent can be written as 'numerator exponent - denominator's exponent'.


For example.

am / an = a ( m - n)


( 26 x 96 ) Divided by  1814 = (18)6 / 1814 = 186 - 14 = 18 - 8 = 1/ (188)


Answer: 18 - 8 = 1/ 188.